# One protein is all you need

### v2

@@ 4.2 Protein Fitness Prediction
We evaluate the models using ProteinGym, the state-of-the-art fitness prediction benchmark (Notin et al., 2024), focusing on its well-established zero-shot setup. Since the zero-shot setup only provides a test set without any data split, we also validate ProteinTTT on independent data. To achieve this, we create a new fitness prediction dataset mined from MaveDB, a public repository containing datasets from Multiplexed Assays of Variant Effect (MAVEs) (Esposito et al., 2019). Following ProteinGym, we measure performance on both datasets using Spearman correlation between predicted and experimental fitness values.
@@ Results.
We observe that ProteinTTT primarily improves performance for proteins with low MSA depth (i.e., the number of available homologous sequences), suggesting that single-sequence customization enhances predictions for proteins with fewer similar sequences in the training data (Table˜A4). The fact that ProteinTTT more effectively improves the performance of smaller ESM2 and SaProt models compared to their larger variants may be a result of the benchmark performance being saturated for larger models, consistent with a recent observation (Notin, 2025). We provide a qualitative example showing how ESM2 (650M) + ProteinTTT significantly improves fitness prediction by capturing residues critical for protein stability (Figure˜A5). We also demonstrate that customization can be combined with evolutionary information from MSA to further boost fitness prediction (Appendix˜B).
@@ 4.3 Protein function prediction
Finally, we demonstrate a proof of concept for customization in the context of protein function prediction. We experiment with two tasks: predicting protein location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (Samusevich et al., 2025). Appendix˜C shows that per-protein customization with ProteinTTT consistently enhances the performance of representative models on both tasks.
@@ 5 Case studies
ProteinTTTcan be readily incorporated into structure, fitness, or function prediction pipelines by adding several lines of code (Appendix˜D). Here, we demonstrate two challenging structure prediction case studies: improving modeling of antibody–antigen loops (Section˜5.1) and expanding known structures of viral proteins (Section˜5.2).
@@ 5.1 Modeling antibody–antigen loops
Accurately predicting structures of antibodies (e.g., human defensive proteins) and antigens (e.g., viral proteins) enables rational design of new therapeutics (Bennett et al., 2025). However, the presence of highly variable loop regions makes modeling of these interactions a long-standing challenge. Here, we show that ProteinTTT substantially improves structure prediction for these loop-formed complementarity-determining regions (CDRs) of antibodies, i.e., the parts that bind antigens, as well as for antigens themselves, on the well-established SAbDab dataset (Dunbar et al., 2014).
We take the structures from SAbDab that are not predicted well by ESMFold (pLDDT < 70) and show that ProteinTTT improves the LDDT score for 115 of 175 antibody CDR substructures (66%) and 487 of 814 antigen chains (60%). As shown in Figure˜4a, ESMFold + ProteinTTT achieves significantly higher average LDDT scores compared to general-purpose ESMFold. Figure˜4b illustrates how ProteinTTT enables accurate prediction of all three CDRs in an antibody chain, providing an improved understanding of its binding interface with the corresponding antigen.
@@ 5.2 Expanding known structures of viral proteins
Predicting the structures of viral proteins is vital for vaccine development, antiviral design, and understanding infection (Bravi, 2024). Nevertheless, it remains challenging due to the high mutation rate, which often leaves viral proteins without close homologs or experimental structures in databases (Kim et al., 2025). Here, we demonstrate that per-protein customized predictions with ESMFold + ProteinTTT improve viral protein structure prediction, substantially expanding the Big Fantastic Virus Database—the repository of viral protein structures (Kim et al., 2025).
Among all the entries in BFVD, predicted with AlphaFold2 through ColabFold (Mirdita et al., 2022) using MSAs constructed from Logan (Chikhi et al., 2024), only 55% have high-quality structure predictions (pLDDT > 70). We apply ESMFold and ESMFold + ProteinTTT to the BFVD entries to expand the database with higher-quality structures. This is achieved by applying all three methods to the specific protein and taking the predicted structure with the highest pLDDT. While ESMFold manages to improve the predicted structure (as measured by pLDDT) for 10% of the BFVD proteins, ESMFold + ProteinTTT leads to an improvement for 19% of the dataset entries, substantially increasing the quality of known viral protein structures (Figure˜5a).
We validate that the improved pLDDT confidence values from ESMFold + ProteinTTT correlate with the quality of the predicted structures, as measured by LDDT against reference AlphaFold2 structures having pLDDT > 90 (Pearson = 0.875; Figure˜A9). Notably, the largest improvements in pLDDT align with the largest improvements in LDDT (Figure˜5b). We find that the benefit of customization saturates with the number of homologs available for a protein, indicating that ProteinTTT is most effective for challenging, out-of-distribution proteins (Figure˜5c). Finally, Figure˜5d–g shows examples where ProteinTTT enables high-confidence structure predictions in cases where general-purpose, uncustomized AlphaFold2 and ESMFold struggle.
@@ Acknowledgments
We thank Milot Mirdita for discussions and feedback about this work. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through projects e-INFRA CZ [ID:90254], ELIXIR [LM2023055], CETOCOEN Excellence CZ.02.1.01/0.0/0.0/17_043/0009632, ESFRI RECETOX RI LM2023069. This work was also supported by the CETOCOEN EXCELLENCE Teaming project supported from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857560. This work was also supported by the European Union (CLARA (No. 101136607), ERC FRONTIER (no. 101097822), ELIAS (No. 101120237)) and the Technology Agency of the Czech Republic under the NCC Programme from the state budget (No. RETEMED TN02000122) and under the NRP from the EU RRF (No. TEREP TN02000122/001N). This work was also supported by the Czech Science Foundation (GA CR) grant 21-11563M and by the European Union’s Horizon Europe program (ERC, TerpenCode, 101170268). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. M.S. acknowledges support by the National Research Foundation of Korea grants (RS-2020-NR049543, RS-2021-NR061659 and RS-2021-NR056571, RS-2024-00396026), Creative-Pioneering Researchers Program and Novo Nordisk Foundation (NNF24SA0092560).
@@
appendix.Aappendix.Bappendix.Cappendix.Dappendix.Esubsection.E.1subsubsection.E.1.1subsubsection.E.1.2subsubsection.E.1.3subsection.E.2subsubsection.E.2.1subsubsection.E.2.2subsubsection.E.2.3subsection.E.3subsubsection.E.3.1subsubsection.E.3.2subsubsection.E.3.3appendix.Fsubsection.F.1subsection.F.2appendix.Gsubsection.G.1subsection.G.2subsection.G.3
@@ Appendix A Justification of customization via perplexity minimization
While the paradigm of test-time customization has been investigated in other domains, the reasons behind its surprising effectiveness are not completely clear (Liu et al., 2021; Zhao et al., 2023). Here, we offer a potential justification for the effectiveness of ProteinTTT by linking it to perplexity minimization.
Perplexity has traditionally been used in natural language processing to evaluate how well models comprehend sentences (Brown, 2020; Chelba et al., 2013). Protein language modeling has adopted this metric to assess how effectively models “understand” amino acid sequences (Hayes et al., 2024; Lin et al., 2023). For bidirectional, random masking language models, which are the focus of this study, we consider the following definition of perplexity [^2] Please note that this is an approximation of perplexity, which is computationally intractable for bidirectional models, and is often referred to as pseudo-perplexity (Lin et al., 2023; Salazar et al., 2019).:
@@
Several studies have shown that lower perplexity on held-out protein sequences (calculated through the self-supervised track $g\circ f$) correlates with better performance on downstream tasks (via the supervised track $h\circ f$), such as predicting protein contacts (Rao et al., 2020), structure (Lin et al., 2023), or fitness (Kantroo et al., 2024). To give an example, we analyze the correlation between perplexity and structure prediction quality (Figure˜A1; see Section˜4.1 for experimental details). A notable correlation suggests that reducing a model’s perplexity on a single target sample $x$ (applied independently to all test samples) can lead to improved predictions on the downstream task (Figure˜3; Figure˜A10).
Since we assume only a single target example $x$, the minimization of the masked language modeling loss $\mathcal{L}(x;\theta)$ (Equation˜2) on this example is directly linked to minimizing the perplexity Perplexity(x)(x) (Equation˜3). For instance, in the case of a single masked position (i.e., $|M|=1$), the loss is equal to the logarithm of perplexity. More generally, it can be shown formally that by minimizing the masked language modeling objective, the model learns to approximate the conditional marginals of the language (of proteins), including the leave-one-out probabilities evaluated in perplexity (Hennigen & Kim, 2023). As a result, applying self-supervised test-time customization on $x$ through $g\circ f$ enhances the representation of the target protein in the backbone $f$, leading to improved downstream performance via the fine-tuning track $h\circ f$.
Appendix B Customization with multiple sequence alignment (MSA)
@@
Since many target proteins may not have homologous sequences (Rao et al., 2021) and finding such homologs may be time-consuming (Lin et al., 2023), the ProteinTTT customization objective (Equation˜2) only assumes a single target sequence for customization. However, we also extend the loss function to the case when a multiple sequence alignment (MSA) is available:
@@
0.4299 ± 0.00099
MSA Transformer (Rao et al., 2021)
0.4319
MSA Transformer + ProteinTTT (Ours)
0.4326 ± 0.00003
[…]
Appendix C Customization for protein function prediction
Protein function prediction is essential for understanding biological processes and guiding bioengineering, but is challenging due to its vague definition and limited data (Yu et al., 2023; Radivojac & et al., 2013; Stärk et al., 2021; Mikhael et al., 2024; Samusevich et al., 2025). While improved structure prediction with ProteinTTT (Section˜4.1) can already enhance function prediction (Song et al., 2024), we also evaluate our customization method directly on two function classification tasks: subcellular localization, predicting protein location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (Christianson, 2017; Samusevich et al., 2025). Using ProteinTTT with EnzymeExplorer (Samusevich et al., 2025) for TPS detection and Light attention (Stärk et al., 2021) for subcellular localization, we achieve consistent performance gains.
@@ Appendix C Customization for protein function prediction
For the terpene substrate classification, we use the largest available dataset of characterized TPS from Samusevich et al. (2025) and reuse the original cross-validation schema. In the case of protein localization prediction, we use a standard DeepLoc dataset (Almagro Armenteros et al., 2017) as a validation set and setHard from Stärk et al. (2021) as the test set.
Given a protein, the goal of function prediction is to correctly classify it into one of the predefined functional annotations. We assess the quality of the TPS substrate prediction using standard multi-label classification metrics used in the EnzymeExplorer paper (Samusevich et al., 2025): mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC). In the case of protein localization prediction, we similarly use the classification metrics from the original paper (Stärk et al., 2021): accuracy, multi-class Matthews correlation coefficient (MCC), and F1-score.
@@ Results.
Table A2: Customization with ProteinTTT improves protein function prediction. For the terpene syntase (TPS) substrate classification task, the metrics are computed on the 512 TPS sequences based on the cross-validation schema of the TPS dataset (Samusevich et al., 2025). Subcellular localization prediction performance is reported for 432 protein sequences from the setHard test set (Stärk et al., 2021). The error bars show standard deviations across five random seeds.
@@
Method
mAP $\uparrow$ AUROC $\uparrow$
EnzymeExplorer
(Samusevich et al., 2025)
0.805 0.948
EnzymeExplorer + ProteinTTT
0.811 ± 0.0011 0.950 ± 0.0002
@@
Method
Accuracy $\uparrow$ MCC $\uparrow$ F1-score $\uparrow$
Light attention
(Stärk et al., 2021)
0.627 0.549 0.618
Light attention + ProteinTTT
0.634 ± 0.004 0.557 ± 0.005 0.627 ± 0.004
@@
Appendix D Implementation details
Infrastructure.
Source code.
We provide a user-friendly and easily extensible PyTorch (Paszke, 2019) implementation of ProteinTTT, available as the proteinttt Python package  [^3] https://github.com/anton-bushuiev/ProteinTTT . We provide two Python code snippets ˜1 and ˜2 to demonstrate the implementation of inference and customization with ProteinTTT, respectively. ˜1 demonstrates how inference with ESMFold can be enhanced with ProteinTTT by adding just a few lines of code to enable customization. Next, ˜2 shows how ProteinTTT can be easily implemented for a PLM of interest by inheriting from the abstract TTTModule class. To integrate ProteinTTT within a model (e.g., ESM2), the user needs to implement methods that define the model’s vocabulary, an interface for predicting logits, and a specification of which modules need to be fine-tuned or remain frozen. The rest, i.e., the test-time training logic itself, is implemented within the unified TTTModule class.
Listing 1: Incorporation of ProteinTTT into an ESMFold structure prediction pipeline using the proteinttt package.

### v3

@@ 4.2 Protein Fitness Prediction
We evaluate the models using ProteinGym, the state-of-the-art fitness prediction benchmark (notin2024proteingym), focusing on its well-established zero-shot setup. Since the zero-shot setup only provides a test set without any data split, we also validate ProteinTTT on independent data. To achieve this, we create a new fitness prediction dataset mined from MaveDB, a public repository of Multiplexed Assays of Variant Effect (MAVEs) (esposito2019mavedb). Following ProteinGym, we report Spearman correlation between predicted and experimental fitness values.
@@ Results.
We observe that ProteinTTT primarily improves performance for proteins with low MSA depth (i.e., the number of available homologous sequences), suggesting that single-sequence customization enhances predictions for proteins with fewer similar sequences in the training data (Table˜A4). The fact that ProteinTTT more effectively improves the performance of smaller ESM2 and SaProt models compared to their larger variants may be a result of the benchmark performance being saturated for larger models, consistent with a recent observation (notin2025scalingwall). We provide a qualitative example showing how ESM2 (650M) + ProteinTTT significantly improves fitness prediction by capturing residues critical for protein stability (Figure˜A6). We also demonstrate that customization can be combined with evolutionary information from MSA to further boost fitness prediction (Appendix˜C).
@@ 4.3 Protein function prediction
Finally, we demonstrate a proof of concept for customization in the context of protein function prediction. We experiment with two tasks: predicting protein location within a cell (stark2021light), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (samusevich2024discovery). Appendix˜D shows that per-protein customization with ProteinTTT consistently enhances the performance of representative models on both tasks.
@@ 5 Case studies
ProteinTTTcan be incorporated into structure, fitness, or function prediction pipelines with a few lines of code (Appendix˜E). Here, we demonstrate two challenging case studies: improving modeling of antibody–antigen loops (Section˜5.1) and expanding known structures of viral proteins (Section˜5.2).
@@ 5.1 Modeling antibody–antigen loops
Accurately predicting structures of antibodies (e.g., human defensive proteins) and antigens (e.g., viral proteins) enables rational design of new therapeutics (bennett2025atomically). However, the presence of highly variable loop regions makes modeling of these interactions a long-standing challenge. Here, we show that ProteinTTT substantially improves structure prediction for these loop-formed complementarity-determining regions (CDRs) of antibodies, i.e., the parts that bind antigens, as well as for antigens themselves, on the well-established SAbDab dataset (dunbar2014sabdab).
We take the structures from SAbDab that are not predicted well by ESMFold (pLDDT < 70) and show that ProteinTTT improves the LDDT score for 115 of 175 antibody CDR substructures (66%) and 487 of 814 antigen chains (60%). As shown in Figure˜4a, ESMFold + ProteinTTT achieves significantly higher average LDDT scores compared to general-purpose ESMFold (paired t-test p-value < 0.05). Figure˜4b illustrates how ProteinTTT enables accurate prediction of all three CDRs in an antibody chain, providing an improved understanding of its binding interface with the corresponding antigen.
@@ 5.2 Expanding known structures of viral proteins
Predicting the structures of viral proteins is vital for vaccine development, antiviral design, and understanding infection (bravi2024development). Nevertheless, it remains challenging due to the high mutation rate, which often leaves viral proteins without close homologs or experimental structures in databases (kim2025bfvd). Here, we demonstrate that per-protein customized predictions with ESMFold + ProteinTTT improve viral protein structure prediction, substantially expanding the Big Fantastic Virus Database—the comprehensive repository of 351,242 viral protein structures (kim2025bfvd).
Among all the entries in BFVD, predicted with AlphaFold2 through ColabFold (mirdita2022colabfold) using MSAs constructed from Logan (chikhi2024logan), only 55% have high-quality structure predictions (pLDDT > 70). We apply ESMFold and ESMFold + ProteinTTT to the BFVD entries to expand the database with higher-quality structures. This is achieved by applying all three methods to the specific protein and taking the predicted structure with the highest pLDDT. While ESMFold manages to improve the predicted structure (as measured by pLDDT) for 10% of the BFVD proteins, ESMFold + ProteinTTT leads to an improvement for 19% of the dataset entries, substantially increasing the quality of known viral protein structures (Figure˜5a).
We validate that the improved pLDDT confidence values from ESMFold + ProteinTTT correlate with the quality of the predicted structures, as measured by LDDT against reference AlphaFold2 structures having pLDDT > 90 (Pearson = 0.875; Figure˜A11). Notably, the largest improvements in pLDDT align with the largest improvements in LDDT (Figure˜5b). We find that the benefit of customization saturates with the number of homologs available for a protein, indicating that ProteinTTT is most effective for challenging, out-of-distribution proteins (Figure˜5c). Finally, Figure˜5d–g shows examples where ProteinTTT enables high-confidence structure predictions in cases where general-purpose, uncustomized AlphaFold2 and ESMFold struggle.
@@ Acknowledgments
We thank Milot Mirdita for discussions and feedback about this work. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through projects e-INFRA CZ [ID:90254], ELIXIR [LM2023055], CETOCOEN Excellence CZ.02.1.01/0.0/0.0/17_043/0009632, ESFRI RECETOX RI LM2023069. This work was also supported by the CETOCOEN EXCELLENCE Teaming project supported from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857560. This work was also supported by the European Union, ERC FRONTIER (No. 101097822), ELIAS (No. 101120237)) and the Technology Agency of the Czech Republic under the NCC Programme from the state budget (No. RETEMED TN02000122) and under the NRP from the EU RRF (No. TEREP TN02000122/001N). This work was also supported by the EU’s Horizon Europe Programme under the Grant agreement No. 101136607 (CLARA Project), and was co-funded by the EU from the Operational Programme Jan Amos Komenský (OP JAK) (project “Center for Artificial Intelligence and Quantum Computing in System Brain Research, reg. no. CZ.02.01.01/00/23_029/0008437). This work was also supported by the Czech Science Foundation (GA CR) grant 21-11563M and by the European Union’s Horizon Europe program (ERC, TerpenCode, 101170268). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. M.S. acknowledges support by the National Research Foundation of Korea grants (RS-2020-NR049543, RS-2021-NR061659 and RS-2021-NR056571, RS-2024-00396026), Creative-Pioneering Researchers Program and Novo Nordisk Foundation (NNF24SA0092560). This work was supported with funding from Google.org via the Google PhD Fellowship (R.B.).
@@
appendix.Aappendix.Bappendix.Cappendix.Dappendix.Eappendix.Fsubsection.F.1subsubsection.F.1.1subsubsection.F.1.2subsubsection.F.1.3subsection.F.2subsubsection.F.2.1subsubsection.F.2.2subsubsection.F.2.3subsection.F.3subsubsection.F.3.1subsubsection.F.3.2subsubsection.F.3.3appendix.Gsubsection.G.1subsection.G.2appendix.Hsubsection.H.1subsection.H.2subsection.H.3appendix.I
@@ Appendix A Justification of customization via perplexity minimization
While the paradigm of test-time customization has been investigated in other domains, the reasons behind its surprising effectiveness are not completely clear (liu2021ttt; zhao2023pitfalls). Here, we offer a potential justification for the effectiveness of ProteinTTT by linking it to perplexity minimization.
Perplexity has traditionally been used in natural language processing to evaluate how well models comprehend sentences (brown2020language; chelba2013one). Protein language modeling has adopted this metric to assess how effectively models “understand” amino acid sequences (hayes2024simulating; lin2023evolutionary). For bidirectional, random masking language models, which are the focus of this study, we consider the following definition of perplexity [^2] Please note that this is an approximation of perplexity, which is computationally intractable for bidirectional models, and is often referred to as pseudo-perplexity (lin2023evolutionary; salazar2019masked).:
@@
Several studies have shown that lower perplexity on held-out protein sequences (calculated through the self-supervised track $g\circ f$) correlates with better performance on downstream tasks (via the supervised track $h\circ f$), such as predicting protein contacts (rao2020transformer), structure (lin2023evolutionary), or fitness (kantroo2024pseudo). To give an example, we analyze the correlation between perplexity and structure prediction quality (Figure˜A1; see Section˜4.1 for experimental details). A notable correlation suggests that reducing a model’s perplexity on a single target sample $x$ (applied independently to all test samples) can lead to improved predictions on the downstream task (Figure˜3; Figure˜A9).
Since we assume only a single target example $x$, the minimization of the masked language modeling loss $\mathcal{L}(x;\theta)$ (Equation˜2) on this example is directly linked to minimizing the perplexity Perplexity(x)(x) (Equation˜3). For instance, in the case of a single masked position (i.e., $|M|=1$), the loss is equal to the logarithm of perplexity. More generally, it can be shown formally that by minimizing the masked language modeling objective, the model learns to approximate the conditional marginals of the language (of proteins), including the leave-one-out probabilities evaluated in perplexity (hennigen2023deriving). As a result, applying self-supervised test-time customization on $x$ through $g\circ f$ enhances the representation of the target protein in the backbone $f$, leading to improved downstream performance via the fine-tuning track $h\circ f$.
Appendix B Customization beyond masked language modeling
In this work, we primarily focus on protein language models pretrained with masked language modeling (MLM), where a fixed proportion of randomly selected tokens (e.g., 15%) are masked for training. To date, MLM has been the dominant paradigm in protein representation learning. Nevertheless, we also provide a proof of concept showing that ProteinTTT can be applied to autoregressive and discrete diffusion–based protein language models, with details provided in the corresponding paragraphs below. Furthermore, in Appendix˜I we discuss how ProteinTTT could be extended beyond protein language models.
Autoregressive customization objective.
To perform single-sequence customization in an autoregressive setting (i.e., customization of ProGen2 (nijkamp2023progen2)), we apply a standard teacher forcing procedure (vaswani2017attention) with a batch size of one. Specifically, each ProteinTTT step optimizes next token prediction across the whole sequence in parallel via the following loss function:
$\mathcal{L}_{\text{AR}}(x;\theta)=\frac{1}{|x|}\sum_{i=1}^{|x|}-\log p(x_{i}\mid x_{<i};\theta),$ (4)
where $x$ denotes a sequence of protein tokens, and $p(x_{i}\mid x_{<i};\theta)\doteq g(f(x_{<i};\theta))_{x_{i}}$ is the probability assigned by the model to the true token $x_{i}$ given all preceding tokens $x_{<i}$. Here, we use the notation consistent with Equation˜2.
Discrete diffusion customization objective.
Recently, discrete diffusion protein language models have emerged as an extension of MLM-based protein language models. Instead of masking a fixed ratio of tokens, discrete diffusion approaches vary the masking ratio during training according to a diffusion schedule (hsieh2025elucidating; wang2024dplm; wang2024diffusion; campbell2024generative; alamdari2023protein). This has been shown to improve representation learning and to enable sequence generation by starting from a fully masked sequence and gradually denoising it (wang2024diffusion).
In this work, we experiment with the DPLM2 Bit-based discrete diffusion model (hsieh2025elucidating) for protein structure prediction. Interestingly, we find that using a standard MLM objective with a fixed 15% masking ratio for customization (Equation˜2) already improves performance. Exploring modifications of the customization objective tailored specifically to discrete diffusion models presents an exciting direction for future work.
Appendix C Customization with multiple sequence alignment (MSA)
Method Avg. Spearman $\uparrow$
ESM2 (lin2023evolutionary) 0.4139
[…]
MSA Transformer (rao2021msa) 0.4319
MSA Transformer + ProteinTTT (Ours) 0.4326 ± 0.00003
@@ Appendix C Customization with multiple sequence alignment (MSA)
Since many target proteins may not have homologous sequences (rao2021msa) and finding such homologs may be time-consuming (lin2023evolutionary), the ProteinTTT customization objective (Equation˜2) only assumes a single target sequence for customization. However, we also extend the loss function to the case when a multiple sequence alignment (MSA) is available:
@@
Appendix D Customization for protein function prediction
Protein function prediction is essential for understanding biological processes and guiding bioengineering, but is challenging due to its vague definition and limited data (yu2023enzyme; radivojac2013large; stark2021light; mikhael2024clipzyme; samusevich2024discovery). While improved structure prediction with ProteinTTT (Section˜4.1) can already enhance function prediction (song2024accurately), we also evaluate our customization method directly on two function classification tasks: subcellular localization, predicting protein location within a cell (stark2021light), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (christianson2017structural; samusevich2024discovery). Using ProteinTTT with EnzymeExplorer (samusevich2024discovery) for TPS detection and Light attention (stark2021light) for subcellular localization, we achieve consistent performance gains.
@@ Appendix D Customization for protein function prediction
For the terpene substrate classification, we use the largest available dataset of characterized TPS from samusevich2024discovery and reuse the original cross-validation schema. In the case of protein localization prediction, we use a standard DeepLoc dataset (almagro2017deeploc) as a validation set and setHard from stark2021light as the test set.
Given a protein, the goal of function prediction is to correctly classify it into one of the predefined functional annotations. We assess the quality of the TPS substrate prediction using standard multi-label classification metrics used in the EnzymeExplorer paper (samusevich2024discovery): mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC). In the case of protein localization prediction, we similarly use the classification metrics from the original paper (stark2021light): accuracy, multi-class Matthews correlation coefficient (MCC), and F1-score.
@@ Results.
Table A2: Customization with ProteinTTT improves protein function prediction. For the terpene syntase (TPS) substrate classification task, the metrics are computed on the 512 TPS sequences based on the cross-validation schema of the TPS dataset (samusevich2024discovery). Subcellular localization prediction performance is reported for 432 protein sequences from the setHard test set (stark2021light). The error bars show standard deviations across five random seeds.
@@
Method mAP $\uparrow$ AUROC $\uparrow$
EnzymeExplorer (samusevich2024discovery) 0.805 0.948
EnzymeExplorer + ProteinTTT (Ours) 0.811 ± 0.0011 0.950 ± 0.0002
@@
Method Accuracy $\uparrow$ MCC $\uparrow$ F1-score $\uparrow$
Light attention (stark2021light) 0.627 0.549 0.618
Light attention + ProteinTTT (Ours) 0.634 ± 0.004 0.557 ± 0.005 0.627 ± 0.004
@@
Appendix E Implementation details
Code snippet 1: Incorporation of ProteinTTT into an ESMFold structure prediction pipeline using the proteinttt package.